Use the definition of Taylor series to find the Taylor series (centered at c) for the function. f(x) = 8x, c=0 f(x) =
Answer #1


* [e=1] Given f(n) = esa, cao f(0) = 8 & f(0) = 0 flow gel {/102=8 f (W) = 8² 8x + (0) = 87 f (x) = 83 8x + (0) = 83 + 1 = 8